# When does bottom-up beat top-down in hierarchical community detection?

### v2

@@ Abstract
Hierarchical clustering of networks consists in finding a tree of communities in order that higher depths of the hierarchy reveal finer-grained community structures. There are two main classes of algorithms for this problem. Divisive (top-down) algorithms recursively partition the nodes into two (or more) communities, until a stopping rule indicates that no further split is needed. In contrast, agglomerative (bottom-up) algorithms first identify the smallest community structure and then repeatedly merge the communities by using a linkage method. In this article, we establish theoretical guarantees for recovering the hierarchy in the community structure of a hierarchical stochastic block model (HSBM) by a bottom-up algorithm. We also establish that this bottom-up algorithm attains the information-theoretic threshold for exact recovery at intermediate depths of the hierarchy. Notably, these recovery conditions are less restrictive compared to those existing for top-down algorithms. This shows that bottom-up algorithms extend the feasible region for achieving exact recovery at intermediate depths. Numerical experiments on both synthetic and real data sets confirm the superiority of bottom-up algorithms over top-down algorithms. In particular, a notable drawback of top-down algorithms is their tendency to produce dendrograms with inversions. These findings contribute to a better understanding of hierarchical clustering techniques and their applications in network analysis.
Keywords: Community detection, hierarchical clustering, linkage, stochastic block model.
@@ 1 Introduction
A system of pairwise interactions among entities can conveniently be represented by a graph, where the system entities are the nodes and the interactions are the edges. Data collected in such form is increasingly abundant in many disciplines, such as sociology, physics, economics, and biology New (18). Finding community structures by grouping nodes with similar connection patterns into clusters is one of the most important statistical analysis tasks on networks For (10); AD (22). Community structures are often hierarchical. For example, in a co-authorship network, we can partition the researchers, based on their primary discipline (such as mathematics, physics, computer science, etc); each of these fields can be further split into specific sub-disciplines. The sub-division of larger communities into smaller ones provides a finer division of the network.
@@
Hierarchical communities can be inferred using top-down approaches, where the process begins by identifying the largest communities at the top of the hierarchy. These communities are then recursively decomposed into smaller sub-communities until a stopping rule indicates no further division is necessary. To identify the communities at the top of the hierarchy, one strategy is to progressively remove edges with the highest edge-betweenness centrality GN (02) or with the lowest edge-clustering coefficient RCC+ (04). Another strategy is to bi-partition the network by using spectral clustering  DHKM (06); BXKS (11); LLB+ (22). Top-down algorithms, however, possess several limitations. First, any clustering errors initially made become locked in and propagate to the next rounds of subdivisions, potentially compromising the accuracy of the predicted hierarchy. Second, the recursive splittings can overlook valuable information by ignoring the edges between communities that have already been separated. As a result, these approaches might not fully capture the interconnections between communities across different depths of the hierarchy.
In contrast, agglomerative (or bottom-up) algorithms take a different approach by constructing the hierarchy from the bottom upwards. These algorithms recursively merge smaller communities to form larger ones. Some bottom-up algorithms, such as the one proposed in BCGH (18); PL (05); CHCL (11), generate a complete dendrogram. A dendrogram is a tree whose leaves are individual nodes, whose branches and internal nodes represent merged clusters, and the lengths of the branches are proportional to the degree of similarity between the merged clusters. Other algorithms, such as greedy schemes maximizing modularity BGLL (08); HSH+ (10), produce partial dendrograms because the leaves are no longer associated with individual nodes but with the bottom communities. Alternatively, initially inferring the bottom flat communities using a graph-clustering algorithm and then merging these communities, through a procedure known as linkage CAKMTM (19), also produces a partial dendrogram. This last approach offers flexibility in building the dendrogram, by selectively incorporating community merges based on specific criteria.
@@
Establishing theoretical guarantees by assessing the performance of HCD algorithms on random graph models is important to compare them. The hierarchical stochastic block model (HSBM) is a general model of random graphs with hierarchical communities. This model defines the hierarchical community structure as a rooted binary tree whose leaves correspond to primitive communities. Each node belongs to a primitive community, and the interactions between two nodes belonging to communities $a$ and $b$ depend only on the lowest common ancestor between $a$ and $b$ on the hierarchical tree.
@@
We first establish that recovering the hierarchy of an HSBM is possible using a bottom-up algorithm when the average degree of the graph is only $\omega(1)$. Earlier studies required stronger degree growth conditions, such as $\Theta(N)$ in BXKS (11), $\omega(N^{1/2}\log^{p}N)$ (with $p=1/2$ in CAKMT (17) and $p=2$ in LTA+ (16)), or $\omega(\log^{1+p}N)$ (with $p=5$ for DHKM (06) and $p=1+\Omega(1)$ for LLB+ (22)).
@@
In the context of hierarchical communities, it is possible to study community recovery at different depths of the hierarchy. We rigorously establish the information-theoretic threshold for the exact recovery of the communities at any intermediate depth of the hierarchy. Exact recovery is the strongest notion of recovery, as it is defined by the ability to fully recover the communities when the number of nodes $N$ tends to infinity. In particular, we establish that exact recovery at a higher depth is more challenging than exact recovery at a lower depth.
We also demonstrate that a bottom-up algorithm using the linkage procedure achieves this threshold. Previous research in this area showed that recursive spectral bi-partitioning, a top-down algorithm, recovers the depths 00 to $\ell$ of the hierarchy, asymptotically exactly if a set of conditions is respected LLB+ (22). Our work reveals that these conditions are stricter than the information-theoretic threshold. As a result, bottom-up algorithms extend the feasible region of exact recovery at intermediate depths, up to the information-theoretic limits.
@@
Finally, in our numerical experiments, we employ a bottom-up algorithm that follows a two-step process. To infer the underlying primitive communities, we first apply a spectral algorithm on the graph’s Bethe-Hessian matrix  SKZ (14); DCT (21). Next, we build the hierarchy using the linkage procedure. To provide a comprehensive comparison, we evaluate this bottom-up algorithm against recursive spectral bi-partitioning Lei (20); LLL (21), as it is the most relevant top-down algorithm whose performance has been theoretically established and numerically validated. Our findings on synthetic data sets demonstrate that the bottom-up algorithm achieves exact recovery at intermediate depths, up to the information-theoretic thresholds, whereas recursive spectral bi-partitioning fails to do so. Furthermore, we show that the dendrogram produced by recursive spectral bi-partitioning can suffer from inversions. Inversions occur when the clustering algorithm incorrectly places a lower-depth cluster above a higher-depth cluster in the dendrogram, thus distorting the true hierarchical structure. Such inversions can lead to misleading interpretations of the hierarchical relationships within the data.
@@
The paper is structured as follows. In Section 2, we describe top-down and bottom-up approaches for hierarchical community detection. The HSBM is defined in Section 3, where we also derive conditions for recovering the hierarchy by the linkage procedure, as well as its robustness to mistakes made in the initial clustering. Next, we investigate the exact recovery at intermediate depths of the hierarchy in Section 4. We discuss these results in light of the existing literature in Section 5. Finally, Section 6 is devoted to the numerical experiments.
@@
We denote by $[N]$ the set $\{1,\cdots,N\}$, by $\operatorname{Ber}(p)$ a Bernoulli random variable with parameter $p$, and by $\mathrm{Multi}(N,\mathcal{S},\pi)$ a multinomial distribution with $N$ trials, a (finite, discrete) set of possible outcomes $\mathcal{S}$, and a vector of probability $\pi$ such that $\pi_{s}$ is the probability of success of $s\in\mathcal{S}$. The Frobenius norm of a matrix is denoted $\|\cdot\|_{F}$.
The Rényi divergence of order $t\in(0,1)$ of a Bernoulli distribution $\operatorname{Ber}(p)$ from another distribution $\operatorname{Ber}(q)$ is defined as $\operatorname{D}_{t}\left(\operatorname{Ber}(p)\|\operatorname{Ber}(q)\right)=\frac{1}{t-1}\log\left((1-p)^{t}(1-q)^{1-t}+p^{t}q^{1-t}\right)$. For $t=1/2$, we write $\operatorname{D}_{1/2}\left(\operatorname{Ber}(p),\operatorname{Ber}(q)\right)=-2\log\left(\sqrt{(1-p)(1-q)}+\sqrt{pq}\right)$ and this quantity is symmetric in $p$ and $q$.
@@ 2 Hierarchical Community Detection
Many networks present a hierarchical community structure. The primitive communities at the bottom of the hierarchy are a collection of subsets $\mathcal{C}=(C_{1},\cdots,C_{K})$ that partition the original node set $V$ into $K$ disjoint sets. These primitive communities are the leaves of a rooted tree $\mathcal{T}$, and this tree defines the hierarchical relationship between the communities. This section reviews two main strategies for hierarchical community detection (HCD).
@@ 2.1 Divisive (top-down) Algorithms
Different choices for the stopping rule or the bi-partitioning algorithms have been explored DHKM (06); BXKS (11); LLB+ (22).
Each recursive splitting of the graph loses some information. Indeed, consider the two clusters $C_{0}$ and $C_{1}$ obtained after the first split. The next step splits $C_{0}$ (resp., $C_{1}$) into two clusters $C_{00}$ and $C_{01}$ (resp., $C_{10}$ and $C_{11}$), based only on the induced sub-graph $G[C_{0}]$ (resp., $G[C_{1}]$). This implies that the clustering of $C_{0}$ does not take into account the edges from $C_{0}$ to $C_{1}$. If the edge densities between $C_{00}$ and $C_{1}$ are different from the edge densities between $C_{01}$ and $C_{1}$, then this valuable information is not used by a top-down algorithm.
Furthermore, it is important to note that the resulting tree is unweighted. Although it is possible to compute similarities between pairs of predicted clusters (for example using the edge density between two clusters), these similarities are not guaranteed to increase with the depth of the dendrogram; inversions can occur. These inversions, as we will emphasize in the numerical section, lead to discrepancies in the hierarchical structure portrayed by the dendrogram.
@@ 2.2 Agglomerative (bottom-up) Algorithms
A first approach is to initially assign each node to a separate community and then merge the clusters that minimize a distance metric New (04); PL (05); CHCL (11); BCGH (18). Although these methods generate a complete dendrogram, determining the depth in the hierarchy where the community structure becomes meaningful is an old yet challenging problem Moj (77); GBW (24).
@@
      merge the two clusters that are the most similar [^1] When measuring the similarity between pairs of clusters by using edge density, it is implicitly assumed that the clustering structure exhibits assortativity. This means that nodes within the same community are more likely to be connected. However, in cases where the communities are disassortative, the focus shifts towards merging clusters that display the lowest similarity, thus reflecting the lack of intra-community connections., and update the similarities between this new cluster and the existing ones;
@@
The abundant literature on graph clustering provides numerous candidate algorithms for the first stage. We note that, in most practical applications, the true number of bottom clusters $K$ is often unknown and needs to be inferred at this stage as well. In this paper, we use the edge density as the similarity measure between pairs of clusters. The edge density between two node sets $V_{1},V_{2}\subset V$ is defined by
$\rho\left(V_{1},V_{2}\right)\ =\ \frac{w\left(V_{1},V_{2}\right)}{|V_{1}|\cdot|V_{2}|}\quad\text{ where }\quad w\left(V_{1},V_{2}\right)\ =\ \sum_{i\in V_{1}}\sum_{j\in V_{2}}A_{ij},$ (2.1)
@@
$\rho\left(\widehat{C}_{k_{1}\cup k_{2}},\widehat{C}_{\ell}\right)\ =\ \frac{\left|\widehat{C}_{k_{1}}\right|}{\left|\widehat{C}_{k_{1}\cup k_{2}}\right|}\rho\left(\widehat{C}_{k_{1}},\widehat{C}_{\ell}\right)+\frac{\left|\widehat{C}_{k_{2}}\right|}{\left|\widehat{C}_{k_{1}\cup k_{2}}\right|}\rho\left(\widehat{C}_{k_{2}},\widehat{C}_{\ell}\right).$ (2.2)
Therefore, the edge density naturally defines an average-linkage procedure for merging the clusters identified through the initial flat clustering. In particular, this procedure guarantees that the resulting dendrogram will be free from inversions, thus ensuring a coherent hierarchical structure MC (17). Algorithm 1 provides a concise summary of this process.
Input: Graph $G=(V,E)$, flat graph-clustering algorithm flat-clustering.
Output: Hierarchical clustering tree $\widehat{\mathcal{T}}$ and its associated bottom clusters $\widehat{C}_{1},\cdots,\widehat{C}_{\widehat{K}}$.
1 Let $\widehat{K}$ be the number of clusters and $\widehat{\mathcal{C}}=\left(\widehat{C}_{1},\cdots,\widehat{C}_{\widehat{K}}\right)$ the clusters inferred by flat-clustering; For all $k,\ell\in\left[\widehat{K}\right]$, compute $\rho\left(\widehat{C}_{k},\widehat{C}_{\ell}\right)$ as in (2.1);
@@
  • Let $\widehat{C}_{k},\widehat{C}_{\ell}\in\operatorname*{arg\,max}\limits_{C\neq C^{\prime}\in\widehat{\mathcal{C}}}\rho\left(C,C^{\prime}\right)$;
  • Let $\widehat{\mathcal{C}}\leftarrow\widehat{\mathcal{C}}\,\backslash\,\left\{\widehat{C}_{k},\widehat{C}_{\ell}\right\}$;
@@
  • Let $\widehat{\mathcal{C}}\leftarrow\widehat{\mathcal{C}}\cup\left\{\widehat{C}_{k\cup\ell}\right\}$.
Return: The bottom communities $\widehat{\mathcal{C}}$ and the tree $\widehat{\mathcal{T}}$ capturing the sequence of nested merges in the while-loop.
Algorithm 1 Bottom-up hierarchical community detection.
@@ 3 Tree Recovery from the Bottom
In this section, we study the asymptotic performance of bottom-up algorithms on a class of random graphs with hierarchical community structures. We define the model in Section 3.1 and, in Section 3.2, we establish the condition for recovering the hierarchical tree from the bottom communities. In Section 3.3, we show that the average-linkage procedure is robust to errors made in the initial clustering step.

### v3

@@ Abstract
Hierarchical community detection consists in finding a tree of communities where deeper levels of the hierarchy reveal finer-grained structures. There are two main classes of algorithms for this task. Divisive (top-down) algorithms recursively partition nodes into smaller communities until a stopping criterion indicates that no further splits are necessary. In contrast, agglomerative (bottom-up) algorithms first identify the smallest community structures and then repeatedly merge the communities by using a linkage method. In this work, we prove that a bottom-up algorithm recovers the hierarchy of a hierarchical stochastic block model (HSBM) when the average degree grows unbounded. We also establish the information-theoretic threshold for exact recovery at intermediate depths of the hierarchy and highlight its significance in understanding the limitations of top-down algorithms. Numerical experiments on both synthetic and real datasets demonstrate the superiority of bottom-up methods. In particular, a notable drawback of top-down algorithms is their tendency to produce dendrograms with inversions. These findings contribute to a better understanding of hierarchical clustering techniques and their applications in network analysis.
Keywords: Community detection; hierarchical clustering; Stochastic Block Model;Agglomerative clustering.
@@ 1 Introduction
A system of pairwise interactions among entities can conveniently be represented by a graph, where the system entities are the nodes and the interactions are the edges. Data collected in such form is increasingly abundant in many disciplines, such as sociology, physics, economics, and biology Newman2018_networksbook. Finding community structures by grouping nodes with similar connection patterns into clusters is one of the most important statistical analysis tasks on networks fortunato2010community; avrachenkov2022statistical. Community structures are often hierarchical. For example, in a co-authorship network, we can partition the researchers, based on their primary discipline (such as mathematics, physics, computer science, etc); each of these fields can be further split into specific sub-disciplines. The sub-division of larger communities into smaller ones provides a finer division of the network.
@@
Hierarchical communities can naturally be inferred using top-down approaches, where the process begins by identifying the largest communities at the top of the hierarchy. These communities are then recursively decomposed into smaller sub-communities until a stopping rule indicates no further division is necessary. To identify the communities at the top of the hierarchy, one strategy is to progressively remove edges with the highest edge-betweenness centrality girvan2002community or with the lowest edge-clustering coefficient radicchi2004defining. Another strategy is to bi-partition the network by using spectral clustering  dasgupta2006spectral; balakrishnan2011noise; li2022hierarchical. Top-down algorithms, however, possess several limitations. First, any clustering errors initially made become locked in and propagate to the next rounds of subdivisions, potentially compromising the accuracy of the predicted hierarchy. Second, the recursive splittings can overlook valuable information by ignoring the edges between communities that have already been separated. As a result, these approaches might not fully capture the interconnections between communities across different depths of the hierarchy.
In contrast, agglomerative (or bottom-up) algorithms take a different approach by constructing the hierarchy from the bottom upwards. These algorithms recursively merge smaller communities to form larger ones. Some bottom-up algorithms, such as the one proposed in bonald2018hierarchical; pons2005computing; chang2011general, generate a complete dendrogram. A dendrogram is a tree whose leaves are individual nodes, whose branches and internal nodes represent merged clusters. The length of every branch measures the similarity of its children.
@@
Comparing different hierarchical community detection (HCD) algorithms can be done by establishing theoretical guarantees of their performance on random graph models. The hierarchical stochastic block model (HSBM) is a general model of random graphs containing hierarchical communities. This model defines the hierarchical community structure as a rooted binary tree whose leaves correspond to primitive communities. Each node belongs to a primitive community, and the interactions between two nodes belonging to communities $a$ and $b$ depend only on the lowest common ancestor between $a$ and $b$ on the tree.
@@
We first establish that recovering the hierarchy of an HSBM is possible using a bottom-up algorithm when the average degree of the graph is only $\omega(1)$. Earlier studies required stronger degree growth conditions, such as $\Theta(N)$ in balakrishnan2011noise, $\omega(N^{1/2}\log^{p}N)$ (with $p=1/2$ in cohen2017hierarchical and $p=2$ in lyzinski2016community), or $\omega(\log^{1+p}N)$ (with $p=5$ for dasgupta2006spectral and $p>1$ fixed for li2022hierarchical). We further discuss these results in Section 5.1.
@@
In the context of hierarchical communities, it is possible to study community recovery at different depths of the hierarchy. We rigorously establish the information-theoretic threshold for the exact recovery of the communities at any intermediate depth of the hierarchy. Exact recovery is the strongest notion of recovery: it is the ability to fully recover the communities when the number of nodes $N$ tends to infinity. In particular, we show that exact recovery at a larger depth is more challenging than exact recovery at a smaller depth. We further discuss the implication of this information-theoretic threshold to the success of top-down algorithms in Section 5.2.
@@
In our numerical experiments, we employ a bottom-up algorithm that follows a two-step process. To infer the underlying primitive communities, we first apply a spectral algorithm on the graph’s Bethe-Hessian matrix saade2014spectral; dall2021unified. Next, we use a bottom-up approach to build the hierarchy. To provide a comprehensive comparison, we evaluate this bottom-up algorithm against recursive spectral bi-partitioning lei2020unified; lei2021consistency, the most relevant top-down algorithm whose performance has been theoretically established and numerically validated. Our findings on synthetic data sets demonstrate that the bottom-up algorithm achieves exact recovery at intermediate depths, up to the information-theoretic thresholds, whereas recursive spectral bi-partitioning fails to do so. Furthermore, we show that the dendrogram produced by recursive spectral bi-partitioning suffers from inversions. Inversions occur when the algorithm incorrectly places a lower-depth cluster above a higher-depth cluster in the dendrogram, thus distorting the true hierarchical structure. Such inversions lead to misleading interpretations of the hierarchical relationships within the data.
@@
The paper is structured as follows. In Section 2, we describe top-down and bottom-up approaches for hierarchical community detection. The HSBM is defined in Section 3, where we also derive conditions for recovering the hierarchy by the linkage procedure. Next, we investigate the exact recovery at intermediate depths of the hierarchy in Section 4. We discuss these results in light of the existing literature in Section 5. Finally, Section 6 is devoted to the numerical experiments.
@@
We denote by $[N]$ the set $\{1,\cdots,N\}$, by $\operatorname{Ber}(p)$ a Bernoulli random variable with parameter $p$. The Frobenius norm of a matrix is denoted $\|\cdot\|_{F}$. The Rényi divergence of order $t\in(0,1)$ of a Bernoulli distribution $\operatorname{Ber}(p)$ from another distribution $\operatorname{Ber}(q)$ is defined as $\operatorname{D}_{t}\left(\operatorname{Ber}(p)\|\operatorname{Ber}(q)\right)=\frac{1}{t-1}\log\left((1-p)^{t}(1-q)^{1-t}+p^{t}q^{1-t}\right)$. For $t=1/2$, we write $\operatorname{D}_{1/2}\left(\operatorname{Ber}(p),\operatorname{Ber}(q)\right)=-2\log\left(\sqrt{(1-p)(1-q)}+\sqrt{pq}\right)$.
@@ 2 Hierarchical Community Detection
Many networks present a hierarchical community structure. The primitive communities at the bottom of the hierarchy are a collection of subsets $\mathcal{C}=(C_{1},\cdots,C_{K})$ that partition the original node set $V$ into $K$ disjoint sets. These primitive communities are the leaves of a rooted tree $\mathcal{T}$, which defines the hierarchical relationship between the communities. This section reviews two main strategies for hierarchical community detection (HCD).
@@ 2.1 Divisive (top-down) Algorithms
Different choices for the stopping rule or the bi-partitioning algorithms have been explored dasgupta2006spectral; balakrishnan2011noise; li2022hierarchical.
Each recursive splitting of the graph loses some information. For example, consider the two clusters $C_{0}$ and $C_{1}$ obtained after the first split. The next step splits $C_{0}$ (resp., $C_{1}$) into two subclusters $C_{00}$ and $C_{01}$ (resp., $C_{10}$ and $C_{11}$), based only on the induced sub-graph $G[C_{0}]$ (resp., $G[C_{1}]$). As a result, the clustering of $C_{0}$ does not take into account the edges from $C_{0}$ to $C_{1}$. If the edge densities between $C_{00}$ and $C_{1}$ are different from the edge densities between $C_{01}$ and $C_{1}$, then this valuable information is not used by a top-down algorithm.
Furthermore, the resulting tree is unweighted. Although it is possible to compute similarities between pairs of predicted clusters (for example using the edge density between two clusters), these similarities are not guaranteed to increase with the depth of the dendrogram; inversions can occur. These inversions, as we will emphasize in the numerical section, lead to discrepancies in the hierarchical structure conveyed by the dendrogram.
@@ 2.2 Agglomerative (bottom-up) Algorithms
A first approach initially assigns each node to a separate community and then merge the clusters that minimize a distance metric newman2004fast; pons2005computing; chang2011general; bonald2018hierarchical. Although these methods generate a complete dendrogram, determining the depth in the hierarchy where the community structure becomes meaningful is an old problem mojena1977hierarchical; gao2022selective.
@@
      merge the two clusters that are the most similar, [^1] When measuring the similarity between pairs of clusters using edge density, an implicit assumption is that the clustering structure exhibits assortativity–meaning that nodes within the same community are more likely to be connected. However, if the community structure is disassortative, clusters with the lowest similarity are merged. and update the similarities between this new cluster and the existing ones;
@@
The abundant literature on graph clustering provides numerous candidate algorithms for the first stage. We note that, in most practical applications, the true number of bottom clusters $K$ is often unknown and needs to be inferred at this stage as well. The second stage is commonly called linkage. Various linkage variants exist, depending on the chosen similarity measures and update rules. In this paper, we adopt the linkage variant used in cohen2017hierarchical; cohen2019hierarchical, employing edge density as the similarity measure between pairs of clusters. The edge density between two node sets $V_{1},V_{2}\subset V$ is defined by
$\rho\left(V_{1},V_{2}\right)\ =\ \frac{w\left(V_{1},V_{2}\right)}{|V_{1}|\cdot|V_{2}|}\quad\text{ where }\quad w\left(V_{1},V_{2}\right)\ =\ \sum_{i\in V_{1}}\sum_{j\in V_{2}}A_{ij}.$ (2.1)
@@
[…]
Therefore, the edge density naturally defines an average-linkage procedure for merging the clusters identified through the initial flat clustering. In particular, this procedure guarantees that the resulting dendrogram will be free from inversions, thus ensuring a coherent hierarchical structure murtagh2017algorithms. Algorithm 1 provides a concise summary of this process.
Input: Graph $G=(V,E)$, partition $\widehat{\mathcal{C}}$ of $V$.
[…]
@@
[…]
@@
[…]
Return: The tree $\widehat{\mathcal{T}}$ capturing the sequence of nested merges in the while-loop.
Algorithm 1 Average-linkage.
@@ 3 Tree Recovery from the Bottom
We study the asymptotic performance of Algorithm 1 on a class of random graphs with hierarchical community structures. We define the model in Section 3.1. We prove that Algorithm 1 recovers the hierarchical tree when the average degree of the graph grows unbounded in Section 3.2. Finally, we study the bounded degree regime in Section 3.3.